Milden Company is a merchandiser that plans to sell 29,000 units during the next quarter at a selling price of $57 per unit. The company also gathered the following cost estimates for the next quarter: Cost Cost of good sold Advertising expense Sales commissions Shipping expense Administrative salaries Insurance expense Depreciation expense Cost Formula $27 per unit sold $177,800 per quarter 6% of sales $75,000 per quarter $4.e0 per unit sold $87,800 per quarter $9,760 per quarter $57,000 per quarter Required: 1. Prepare a contribution format income statement for the next quarter. 2. Prepare a traditional format income statement for the next quarter.
Prepare a contribution format income statement for the next quarter. Milden Company Contribution Format Income Statement For the Next Quarter Variable expenses: Total variable expenses Contribution margin Fixed expenses Total fixed expenses
Prepare a traditional format income statement for the next.guarter. Milden Company Traditional Format Income Statement For the Next Quarter Selling and administrative expenses: Total selling and administrative expenses < Required! Required 2 >
Answer #1

Milden Company

Contribution margin Income statement

For the Next Quarter

Sales

$ 16,53,000.00

Less: Variable costs

Variable Cost of goods sold

$   7,83,000.00

Variable selling expenses

$   1,16,000.00

Variable selling commission

$       99,180.00

Total Variable cost

$   9,98,180.00

Contribution margin

$   6,54,820.00

Less: Fixed Costs

Advertising expense

$   1,77,000.00

Shipping expenses

$       75,000.00

Administrative salaries

$       87,000.00

Insurance expense

$         9,700.00

Depreciation expense

$       57,000.00

Total Fixed Costs

$   4,05,700.00

Net Income

$   2,49,120.00

Milden Company

Traditional Format Income Statement

For the Next Quarter

Sales

$ 16,53,000.00

Cost of goods sold

$   7,83,000.00

Gross margin

$   8,70,000.00

Selling and administrative expenses:

Advertising expense

$1,77,000.00

Sales commission

$   99,180.00

Shipping expense

$1,91,000.00

Administrative salaries

$   87,000.00

Insurance expense

$    9,700.00

Depreciation expense

$   57,000.00

Total selling and administrative expenses

$   6,20,880.00

Net operating income

$     2,49,120.00

Working

Milden Company

Traditional Format Income Statement

For the Next Quarter

Sales

=29000*57

Cost of goods sold

=29000*27

Selling and administrative expenses:

Advertising expense

177000

Sales commission

=1653000*6%

Shipping expense

=+(29000*4)+75000

Administrative salaries

87000

Insurance expense

9700

Depreciation expense

57000

Milden Company

Contribution margin Income statement

For the Next Quarter

Sales

=29000*57

Less: Variable costs

Variable Cost of goods sold

=29000*27

Variable selling expenses

=29000*4

Variable selling commission

=1653000*6%

Less: Fixed Costs

Advertising expense

177000

Shipping expenses

75000

Administrative salaries

87000

Insurance expense

9700

Depreciation expense

63000